# middlelinecolor=black, middlelinewidth=0.5pt, backgroundcolor=gray!10, innerleftmargin=4pt, innerrightmargin=4pt, innertopmargin=4pt, innerbottommargin=4pt, roundcorner=4pt

### v1

@@ Abstract
We propose ClassroomKD, a novel multi-mentor knowledge distillation framework inspired by classroom environments to enhance knowledge transfer between student and multiple mentors. Unlike traditional methods that rely on fixed mentor-student relationships, our framework dynamically selects and adapts the teaching strategies of diverse mentors based on their effectiveness for each data sample. ClassroomKD comprises two main modules: the Knowledge Filtering (KF) Module and the Mentoring Module. The KF Module dynamically ranks mentors based on their performance for each input, activating only high-quality mentors to minimize error accumulation and prevent information loss. The Mentoring Module adjusts the distillation strategy by tuning each mentor’s influence according to the performance gap between the student and mentors, effectively modulating the learning pace. Extensive experiments on image classification (CIFAR-100 and ImageNet) and 2D human pose estimation (COCO Keypoints and MPII Human Pose) demonstrate that ClassroomKD significantly outperforms existing knowledge distillation methods. Our results highlight that a dynamic and adaptive approach to mentor selection and guidance leads to more effective knowledge transfer, paving the way for enhanced model performance through distillation.
@@ 1 Introduction
Knowledge distillation (KD) (Hinton et al., 2015) is a widely adopted model compression technique in deep learning, where a smaller, more efficient student model learns to replicate the behavior of a larger, more complex teacher model. While traditional KD methods typically employ a single teacher, multi-teacher (or multi-mentor) distillation has been proposed to further enhance performance by leveraging an ensemble of teachers (You et al., 2017). This setup is expected to provide richer and more diverse knowledge, improving the student’s generalization and robustness. We use the term mentor to describe all networks involved in teaching the student, regardless of their size or role.
Figure 1: (a) DML: Peer models learn from each other without a hierarchical teacher structure. (b) TAKD: A sequential mentor-student hierarchy with large-to-small knowledge transfer. (c) DGKD: Each mentor teaches all smaller models. (d) ClassroomKD: Our proposed method dynamically selects mentors for each data sample based on the current input and ranks them using the Knowledge Filtering Module. (e) Adaptive Mentoring: The Mentoring Module adjusts teaching strategies according to dynamic rankings, ensuring optimal knowledge transfer.
@@
Large Capacity Gap: Employing multiple large mentors can create a substantial capacity gap between the collective representation power of the mentors and that of the student. This gap can hinder the student’s ability to effectively mimic the combined knowledge of the mentors, leading to suboptimal learning outcomes. To bridge this gap, some works (Mirzadeh et al., 2019; Son et al., 2021) have introduced intermediate-sized mentors alongside a large teacher. However, smaller mentors may be less effective, potentially introducing additional errors into the student’s knowledge.
Error Accumulation: The lower performance of smaller mentors can contribute to cumulative errors in the distillation process. This is particularly problematic in sequential distillation frameworks like TAKD (Figure 1(b)), where each mentor teaches only the subsequent smaller model. Such setups can lead to an "error avalanche," where inaccuracies from lower-performing mentors degrade the student’s performance (Son et al., 2021). Although DGKD (Figure 1(c)) attempts to mitigate this by allowing each mentor to teach all smaller models and randomly dropping some mentors during training, these strategies can result in valuable information loss and reduced learning efficiency.
Lack of Dynamic Adaptation: The performance gap between the student and its mentors is not static; it evolves throughout training. Current methods do not adequately address these dynamic scenarios, limiting the effectiveness of multi-mentor distillation (Hao et al., 2024). Without an adaptive strategy, the potential benefits of multi-mentor distillation are not fully realized.
Observing that (1) a mentor’s performance varies across different data samples, (2) each mentor possesses distinct teaching capabilities due to varying capacity gaps, and (3) the performance gap evolves during training, we propose ClassroomKD (Figure 1(d)), a novel multi-mentor distillation framework inspired by classroom dynamics (see Appendix E). Our method introduces two key modules (Figure 1(e)) designed to address the following questions:
@@
We introduce the Knowledge Filtering Module to intelligently select mentors. This module dynamically ranks all mentors based on their performance for each input, activating only those with sufficient performance. A mentor is deemed effective and activated if its predictions are accurate and more confident than the student’s. This minimizes error accumulation and information loss.
@@
Our Mentoring Module addresses this by tuning the teaching strategy based on the performance gap between the student and each active mentor. Specifically, we adjust each mentor’s distillation temperature to control the teaching pace, allowing the student to appropriately weigh information received from each mentor before integrating it into its own knowledge.
@@ 2.1 Knowledge Distillation Approaches
Knowledge distillation (KD) (Hinton et al., 2015) is a widely adopted technique for compressing deep neural networks, where a smaller student model learns from a larger teacher model by minimizing the distance between their output probability distributions, or soft labels. Traditional KD methods primarily focus on logit-based distillation, where the student learns directly from the teacher’s output logits. Notable methods include PKT (Passalis & Tefas, 2018), which employs probabilistic knowledge transfer, FT (Kim et al., 2018), which transfers factorized feature representations, and AB (Heo et al., 2019), which leverages activation boundaries formed by hidden neurons.
Feature-based distillation methods transfer knowledge by aligning intermediate representations between the teacher and student. FitNets (Adriana et al., 2015) introduced this approach using intermediate feature maps for training. Later methods like AT (Zagoruyko & Komodakis, 2016), VID (Ahn et al., 2019), and CRD (Tian et al., 2020) enhance knowledge transfer by matching attention maps, utilizing variational information distillation, and employing contrastive learning, respectively.
Relation-based methods focus on preserving the structural relationships within the teacher’s feature maps. RKD (Park et al., 2019) maintains data point structures through relational knowledge distillation, while SP (Tung & Mori, 2019) and SRRL (Yang et al., 2021) optimize for similarity-preserving objectives. DIST (Huang et al., 2022) addresses large capacity gaps by applying a correlation-based loss to maintain both inter-class and intra-class relationships, enhancing distillation efficiency.
Recent approaches have explored more specialized distillation techniques. WSLD (Zhou et al., 2021) introduces weighted soft labels to balance bias-variance trade-offs, while One-to-All Spatial Matching KD (Lin et al., 2022) focuses on spatial matching techniques. OFA (Hao et al., 2024) optimizes feature-based KD by projecting features onto the logit space, significantly improving performance for heterogeneous models. To enhance distillation effectiveness, several methods have incorporated adaptive strategies. CTKD (Li et al., 2023) dynamically adjusts the temperature during training to gradually increase learning difficulty, and DTKD (Wei & Bai, 2024) employs real-time temperature scaling to improve knowledge transfer efficiency.
@@ 2.2 Multi-Teacher Knowledge Distillation
Multi-teacher distillation methods aim to further enhance student performance by leveraging an ensemble of mentors (You et al., 2017).
Online knowledge distillation has been particularly successful in this context. Deep Mutual Learning (DML) (Zhang et al., 2018) introduces a framework where multiple peer models learn from each other simultaneously during training, fostering collaborative learning among smaller networks and outperforming traditional one-way (offline) distillation. Other online methods include ONE (Zhu et al., 2018), OKDDip (Chen et al., 2020), and FFM (Li et al., 2020), which often outperform offline methods. Online distillation has also been extended to pose estimation tasks (Li et al., 2021). SHAKE (Li & Jin, 2022) proposed using proxy teachers with shadow heads to use the benefits of online distillation in offline settings.
To address the capacity gap in multi-teacher setups, Teacher-Assistant KD (TAKD) (Mirzadeh et al., 2019) employs intermediate-sized teacher assistants (TAs) to bridge the gap between the largest teacher and the student. However, sequential distillation through TAs can result in an "error avalanche," where errors propagate at each step, reducing final performance. Adaptive Ensemble Knowledge Distillation (AEKD) (Du et al., 2020) mitigates this issue by using an adaptive dynamic weighting strategy to reduce error propagation in the gradient space. Densely Guided KD (DGKD) (Son et al., 2021) further improves upon these methods by guiding each TA with both larger TAs and the main teacher, enabling a more gradual and effective transfer of knowledge. Additionally, DGKD introduces a strategy of randomly dropping mentors during training to expose the student to diverse learning sources, enhancing overall learning robustness.
@@ 3.1 Knowledge Filtering Module
$\bm{p}^{m}_{\text{gt}}$ $=\bm{p}^{m}[\bm{y}]$ (3)
@@
This ensures the student learns from high-quality sources by selecting mentors based on their performance ranks. This selective approach prevents error accumulation as only mentors outperforming the student can teach it, avoiding the propagation of errors from less effective mentors. Additionally, it avoids information loss by consistently selecting the best-performing mentors, unlike random mentor-dropping strategies (Son et al., 2021).
@@ 3.2 Mentoring Module
$\mathcal{L}=\alpha\mathcal{L}_{\mathrm{task}}(\hat{\bm{y}}^{s},\bm{y})+\beta\sum_{m\in\mathbb{M}^{\prime}}\gamma^{m}\mathcal{L}_{\mathrm{distill}}(\hat{\bm{y}}^{m},\hat{\bm{y}}^{s};\tau^{m})$ (10)
@@ 4 Experiments
This section presents our experiments to evaluate the effectiveness of ClassroomKD using different datasets. We primarily use CIFAR-100 (Krizhevsky et al., 2009) classification for detailed comparisons with state-of-the-art single and multiple-teacher distillation methods. This also includes online approaches using multiple mentors. In addition, we also report results on ImageNet (Deng et al., 2009) classification and human pose estimation using the COCO Keypoints (Lin et al., 2014) and MPII Human Pose (Andriluka et al., 2014) datasets. Our results show that ClassroomKD outperforms existing methods under various settings, highlighting the robustness and adaptability of our method.
Implementation Details. For CIFAR-100, we train for 240 epochs with a batch size of 64, a learning rate of 0.05 decayed by 10% every 30 epochs, and a 120-epoch warm-up phase. We use SGD with 0.9 momentum and $5\times 10^{-4}$ weight decay. The temperature $\tau$ is set to 12 via grid search (Figure 3). For ImageNet, models are trained for 100 epochs with $\tau=8$. For pose datasets, models are trained for 210 epochs with $\tau=4$. The scaling factor $\lambda$ is $n+1$ for all experiments, where $n$ is the number of peers. We used $\beta=1.0$ for classification and $\beta=2.5$ for pose estimation. We follow standard training protocols, with mentors pre-trained and kept frozen.
@@
4.1 Results
CIFAR-100 Classification: Table 1 compares the performance of ClassroomKD to single-teacher distillation methods on the CIFAR-100 dataset. We evaluate various teacher-student pairings using both homogeneous and heterogeneous architectures. ClassroomKD, which is logit-based, performs better than logit-based methods in a majority of cases, as well as most feature-based and relation-based methods. In particular, when comparing with the recent state-of-the-art CTKD (Li et al., 2023) and DTKD (Wei & Bai, 2024) methods, both of which use adaptive temperatures, ClassroomKD is better than the former in all cases and the latter in 75% of the cases.
@@ 4.1 Results
FitNets (Adriana et al., 2015) 68.99 71.06 69.21 73.54 64.14 73.54 73.73 63.16 78.87
AT (Zagoruyko & Komodakis, 2016) 70.22 72.31 70.55 73.62 59.40 72.73 73.32 - -
VID (Ahn et al., 2019) 70.16 72.61 70.38 73.96 - 73.40 73.61 67.57 -
CRD (Tian et al., 2020) 71.46 73.48 71.16 73.94 69.73 75.65 76.05 69.11 77.63
SimKD (Chen et al., 2022) - - - 74.93 - 77.49 - - -
SMKD (Lin et al., 2022) 71.70 74.05 71.59 74.39 - - - - -
RKD (Park et al., 2019) 69.25 71.82 69.61 73.72 64.52 73.21 72.21 64.43 74.11
SP (Tung & Mori, 2019) 70.04 72.69 69.67 73.44 66.30 74.56 74.52 - -
SRRL (Yang et al., 2021) 71.51 73.80 - 73.23 69.34 75.66 76.61 - -
DIST (Huang et al., 2022) - - 71.75 - - 77.35 - 68.66 77.75
KD (Hinton et al., 2015) 70.67 73.08 70.66 72.98 67.37 74.45 74.83 67.35 78.74
PKT (Passalis & Tefas, 2018) 70.25 72.61 70.34 73.37 - 74.69 73.89 66.52 -
FT (Kim et al., 2018) 70.22 72.37 69.84 73.42 - 72.50 72.03 - -
AB (Heo et al., 2019) 69.53 70.98 69.47 74.27 - 74.31 73.34 - -
WSLD (Zhou et al., 2021) 72.19 74.12 72.15 - - 75.93 76.21 - -
CTKD (Li et al., 2023) 70.99 73.52 71.19 73.52 68.46 75.31 75.78 68.47 -
DTKD (Wei & Bai, 2024) - 74.07 72.05 74.12 69.01 76.19 76.29 69.10 -
OFA (Hao et al., 2024) - - - - - - - - 80.54
Ours (2024) 72.45 74.60 72.65 74.51 69.84 76.52 75.05 70.15 80.32
Table 2: Comparison with multi-teacher distillation methods. Best and second-best values in offline methods are bold and underlined, respectively, and overall best in blue. *:trained on (coco+aic)
(a) Results on CIFAR-100 classification. We report top-1 accuracy (%). KD methods are grouped by online and offline. ClassroomKD is offline.
Method Same Architectures Mixed Architectures

### v2

# Classroom-Inspired Multi-Mentor Distillation with Adaptive Learning Strategies

@@ Abstract
We propose ClassroomKD, a novel multi-mentor knowledge distillation framework inspired by classroom environments to enhance knowledge transfer between the student and multiple mentors with different knowledge levels. Unlike traditional methods that rely on fixed mentor-student relationships, our framework dynamically selects and adapts the teaching strategies of diverse mentors based on their effectiveness for each data sample. ClassroomKD comprises two main modules: the Knowledge Filtering (KF) module and the Mentoring module. The KF Module dynamically ranks mentors based on their performance for each input, activating only high-quality mentors to minimize error accumulation and prevent information loss. The Mentoring Module adjusts the distillation strategy by tuning each mentor’s influence according to the dynamic performance gap between the student and mentors, effectively modulating the learning pace. Extensive experiments on image classification (CIFAR-100 and ImageNet) and 2D human pose estimation (COCO Keypoints and MPII Human Pose) demonstrate that ClassroomKD outperforms existing knowledge distillation methods for different network architectures. Our results highlight that a dynamic and adaptive approach to mentor selection and guidance leads to more effective knowledge transfer, paving the way for enhanced model performance through distillation.
keywords:
knowledge distillation, multi-mentors, lifelong learning, image classification, pose estimation, classroom learning
@@ 1 Introduction
Knowledge distillation (KD) [1] is a widely adopted model compression technique in deep learning, where a smaller, more efficient student model learns to replicate the behavior of a larger, more complex teacher model. While traditional KD methods [1][2][3] typically employ a single teacher, multi-teacher (or multi-mentor) distillation has been proposed to further enhance performance by leveraging an ensemble of teachers [4]. This setup is expected to provide richer and more diverse knowledge, improving the student’s generalization and robustness. We use the term mentor to describe all networks involved in teaching the student, regardless of their size or role.
Figure 1: (a) DML: Peer models learn from each other without a hierarchical teacher structure. (b) TAKD: A sequential mentor-student hierarchy with large-to-small knowledge transfer. (c) DGKD: Each mentor teaches all smaller models. (d) ClassroomKD: Our proposed method dynamically selects mentors for each data sample based on the current input and ranks them using the Knowledge Filtering Module. (e) Adaptive Mentoring: The Mentoring Module adjusts teaching strategies of each active mentor according to dynamic rankings, ensuring optimal knowledge transfer.
@@
Large Capacity Gap: Employing multiple large mentors can create a substantial capacity gap between the collective representation power of the mentors and that of the student. This gap can hinder the student’s ability to effectively mimic the combined knowledge of the mentors, leading to suboptimal learning outcomes. To bridge this gap, some works [5, 6] have introduced intermediate-sized mentors alongside a large teacher. However, smaller mentors may be less effective, potentially introducing additional errors into the student’s knowledge.
Error Accumulation: The lower performance of smaller mentors can contribute to cumulative errors in the distillation process. This is particularly problematic in sequential distillation frameworks like TAKD (Figure 1(b)), where each mentor teaches only the subsequent smaller model. Such setups can lead to an "error avalanche," where inaccuracies from lower-performing mentors degrade the student’s performance [6]. Although DGKD (Figure 1(c)) attempts to mitigate this by allowing each mentor to teach all smaller models and randomly dropping some mentors during training, these strategies can result in valuable information loss and reduced learning efficiency.
Lack of Dynamic Adaptation: The performance gap between the student and its mentors is not static; it evolves throughout training (as visualized in Appendix A.3 Figures 9-11). Current methods do not adequately address these dynamic scenarios, limiting the effectiveness of multi-mentor distillation [7]. Without an adaptive strategy, the potential benefits of multi-mentor distillation are not fully realized.
Observing that (1) a mentor’s performance varies across different data samples, (2) each mentor possesses distinct teaching capabilities due to varying capacity gaps, and (3) the performance gap evolves during training, we draw parallels between knowledge distillation and Vygotsky’s Zone of Proximal Development (ZPD) (1978). His theory emphasizes learning with a More Knowledgeable Other (MKO) and the need for scaffolded support.
We propose ClassroomKD (Figure 1(d)), a novel multi-mentor distillation framework inspired by classroom dynamics (see Appendix F). Our method introduces two key modules (Figure 1(e)) designed to address the following questions:
@@
We introduce the Knowledge Filtering Module to intelligently select mentors(or the MKOs). This module dynamically ranks all mentors based on their performance for each input, activating only those with sufficient performance. A mentor is deemed effective and activated if its predictions are accurate and more confident than the student’s. This minimizes error accumulation and information loss.
@@
Our Mentoring Module addresses this by tuning the teaching strategy based on the performance gap between the student and each active mentor. Specifically, we adjust each mentor’s distillation temperature to control the teaching pace(scaffolding), allowing the student to appropriately weigh information received from each mentor before integrating it into its own knowledge.
@@ 2.1 Knowledge Distillation Approaches
Knowledge distillation (KD) [1] is a widely adopted technique for compressing deep neural networks, where a smaller student model learns from a larger teacher model by minimizing the distance between their output probability distributions, or soft labels. Traditional KD methods primarily focus on logit-based distillation, where the student learns directly from the teacher’s output logits. Notable methods include PKT [8], which employs probabilistic knowledge transfer, FT [9], which transfers factorized feature representations, and AB [10], which leverages activation boundaries formed by hidden neurons.
Feature-based distillation methods transfer knowledge by aligning intermediate representations between the teacher and student. FitNets [3] introduced this approach using intermediate feature maps for training. Later methods like AT [2], VID [11], and CRD [12] enhance knowledge transfer by matching attention maps, utilizing variational information distillation, and employing contrastive learning, respectively.
Relation-based methods focus on preserving the structural relationships within the teacher’s feature maps. RKD [13] maintains data point structures through relational knowledge distillation, while SP [14] and SRRL [15] optimize for similarity-preserving objectives. DIST [16] addresses large capacity gaps by applying a correlation-based loss to maintain both inter-class and intra-class relationships, enhancing distillation efficiency.
Recent approaches have explored more specialized distillation techniques. WSLD [17] introduces weighted soft labels to balance bias-variance trade-offs, while One-to-All Spatial Matching KD [18] focuses on spatial matching techniques. OFA [7] optimizes feature-based KD by projecting features onto the logit space, significantly improving performance for heterogeneous models. To enhance distillation effectiveness, several methods have incorporated adaptive strategies. CTKD [19] dynamically learns the temperature during training to gradually increase learning difficulty, and DTKD [20] employs real-time temperature scaling to improve knowledge transfer efficiency.
@@ 2.2 Multi-Teacher Knowledge Distillation
Multi-teacher distillation methods aim to further enhance student performance by leveraging an ensemble of mentors [4].
Online knowledge distillation has been particularly successful in this context. Deep Mutual Learning (DML) [21] introduces a framework where multiple peer models learn from each other simultaneously during training, fostering collaborative learning among smaller networks and outperforming traditional one-way (offline) distillation. Other online methods include ONE [22], OKDDip [23], and FFM [24], which often outperform offline methods. Online distillation has also been extended to pose estimation tasks [25]. SHAKE [26] proposed using proxy teachers with shadow heads to use the benefits of online distillation in offline settings.
To address the capacity gap in multi-teacher setups, Teacher-Assistant KD (TAKD) [5] employs intermediate-sized teacher assistants (TAs) to bridge the gap between the largest teacher and the student. However, sequential distillation through TAs can result in an "error avalanche", where errors propagate at each step, reducing final performance. Adaptive Ensemble Knowledge Distillation (AEKD) [27] mitigates this issue by using an adaptive dynamic weighting strategy to reduce error propagation in the gradient space. Densely Guided KD (DGKD) [6] further improves upon these methods by guiding each TA with both larger TAs and the main teacher, enabling a more gradual and effective transfer of knowledge. Additionally, DGKD introduces a strategy of randomly dropping mentors during training to expose the student to diverse learning sources, enhancing overall learning robustness.
@@ 3.1 Knowledge Filtering Module
$\bm{p}^{m}_{\text{gt}}$ $=\bm{p}^{m}[\bm{y}]=1/(\text{exp}(\mathrm{CELoss}(\hat{\bm{y}}^{m},\text{targets}))$ (3)
@@
This ensures the student learns from high-quality sources by selecting mentors based on their performance ranks. This selective approach prevents error accumulation as only mentors outperforming the student can teach it, avoiding the propagation of errors from less effective mentors. Additionally, it avoids information loss by consistently selecting the best-performing mentors, unlike random mentor-dropping strategies [6].
@@ 3.2 Mentoring Module
$\mathcal{L}_{\mathrm{classroom}}$ $=\alpha\mathcal{L}_{\mathrm{task}}(\hat{\bm{y}}^{s},\bm{y})+\beta\sum_{m\in\mathbb{M}^{\prime}}\gamma^{m}\mathcal{L}_{\mathrm{distill}}(\hat{\bm{y}}^{m},\hat{\bm{y}}^{s};\tau^{m})$ (10)
$\mathcal{L}$ $=\delta(\mathcal{L}_{\mathrm{task}}(\hat{\bm{y}}^{s},\bm{y})+\mathcal{L}_{\mathrm{distill}}(\hat{\bm{y}}^{t},\hat{\bm{y}}^{s};\tau^{t}=1))+\mathcal{L}_{\mathrm{classroom}}$ (11)
@@ 4 Experiments
This section presents our experiments to evaluate the effectiveness of ClassroomKD using different datasets. We primarily use CIFAR-100 [28] classification for detailed comparisons with state-of-the-art single and multiple-teacher distillation methods. This also includes online approaches using multiple mentors. In addition, we also report results on ImageNet [29] classification and human pose estimation using the COCO Keypoints [30] and MPII Human Pose [31] datasets. Our results show that ClassroomKD outperforms existing methods under various settings, highlighting the robustness and adaptability of our method.
Implementation Details. For CIFAR-100, we train for 240 epochs with a batch size of 64, a learning rate of 0.05 decayed by 10% every 30 epochs, and a 120-epoch warm-up phase. We use SGD with 0.9 momentum and $5\times 10^{-4}$ weight decay. The temperature $\tau$ is set to 12 via grid search (Figure 3). For ImageNet, models are trained for 100 epochs with $\tau=4$. For pose datasets, models are trained for 210 epochs with $\tau=4$. The scaling factor $\lambda$ is $n+1$ for all experiments, where $n$ is the number of peers. We used $\beta=1.0$ for classification and $\beta=2.5$ for pose estimation. Furthermore, we set $\delta=0$ for CIFAR100 dataset and $\delta=1$ for the large-scale Imagenet dataset.
@@
We follow standard training protocols, with mentors pre-trained and kept frozen.
4.1 CIFAR-100 Classification
We begin by comparing ClassroomKD against single-teacher distillation methods in Table 1. The table includes a variety of teacher–student pairs, covering both homogeneous (e.g., ResNet 110 → ResNet 20) and heterogeneous (e.g., VGG 13 → MobileNetV2) architectures.
@@ 4.1 CIFAR-100 Classification
FitNets [3] 68.99 71.06 69.21 73.54 64.14 73.54 73.73 63.16 78.87
AT [2] 70.22 72.31 70.55 73.62 59.40 72.73 73.32 - -
VID [11] 70.16 72.61 70.38 73.96 - 73.40 73.61 67.57 -
CRD [12] 71.46 73.48 71.16 73.94 69.73 75.65 76.05 69.11 77.63
SimKD [32] - - - 74.93 - 77.49 - - -
SMKD [18] 71.70 74.05 71.59 74.39 - - - - -
RKD [13] 69.25 71.82 69.61 73.72 64.52 73.21 72.21 64.43 74.11
SP [14] 70.04 72.69 69.67 73.44 66.30 74.56 74.52 - -
SRRL [15] 71.51 73.80 - 73.23 69.34 75.66 76.61 - -
DIST [16] - - 71.75 - - 77.35 - 68.66 77.75
KD [1] 70.67 73.08 70.66 72.98 67.37 74.45 74.83 67.35 78.74
PKT [8] 70.25 72.61 70.34 73.37 - 74.69 73.89 66.52 -
FT [9] 70.22 72.37 69.84 73.42 - 72.50 72.03 - -
AB [10] 69.53 70.98 69.47 74.27 - 74.31 73.34 - -
WSLD [17] 72.19 74.12 72.15 - - 75.93 76.21 - -
CTKD [19] 70.99 73.52 71.19 73.52 68.46 75.31 75.78 68.47 -
DTKD [20] - 74.07 72.05 74.12 69.01 76.19 76.29 69.10 -
OFA [7] - - - - - - - - 80.54
Ours 72.06 74.71 72.13 75.29 70.26 76.74 75.81 70.23 80.32
Overall Improvements.
ClassroomKD consistently outperforms baseline logit-based methods, as well as many feature-based and relation-based methods. Notably, our method competes favorably with recent approaches that employ adaptive temperature scaling, such as CTKD [19] and DTKD [20]. These gains suggest that our combination of selective mentor activation and dynamically adjusted temperatures more robustly handles the evolving capacity gap than strategies that only tune temperature globally.
Capacity Gap Mitigation.
In large-teacher/small-student pairings (e.g., ResNet-110 → ResNet-20, VGG-13 → MobileNetV2), the capacity gap is significant. ClassroomKD explicitly tackles this by filtering out under-performing mentors in a data-dependent way, reducing “noisy guidance.” This proves especially helpful in preventing a performance plateau observed in many other KD methods when the teacher is much larger.
4.2 CIFAR-100 Classification with Multiple Mentors
We next evaluate the multi-mentor scenario in Table [2(a)], where each classroom includes a single large teacher and several intermediate-capacity peers (details in Appendix D). We compare both online frameworks (e.g., DML [21] and SHAKE [26] and offline ones (e.g., AEKD [27], DGKD [6]).
Defining a Simple Baseline.
Following SOTA methods [26, 33], we use AVER as the simplest baseline in our multi-mentor comparisons. This is a direct counterpart of KD in single-teacher experiments and is defined as:
$\mathcal{L}_{\mathrm{AVER}}=\mathcal{L}_{\mathrm{task}}(\hat{\bm{y}}^{s},\bm{y})+\sum_{m\in\mathbb{M}}\mathcal{L}_{\mathrm{distill}}(\hat{\bm{y}}^{m},\hat{\bm{y}}^{s};\tau)$ (12)
Each teacher is weighted equally without any ranking or temperature adaption; the student naively attempts to learn the aggregate of all teachers’ knowledge.
Table 2: Comparison with multi-teacher distillation methods.
(a) Results on CIFAR-100 classification. We report top-1 accuracy (%). KD methods are grouped by online and offline. ClassroomKD is offline. Best and second-best values in offline methods are bold and underlined, respectively, and overall best in blue. Complete classroom configurations with details about the peers are provided in the appendix.
Method Same Archs Mixed Archs